a man invested a total of $15000 in two venture..in one he made profit of 8% and in other 4% loss..his net income from the two venture is 240$..how much did he invested in each venture?
Answer by checkley77(12844)

.08X-.04(15,000-X)=240
.08X-600+.04X=240
.12X=240+600
.12X=840
X=840/.12
X=%7,000 INVESTED @ 8%.
12,000=7,000=$8,000 INVESTED @ 4%.
PROOF:
.08*7,000-.04(8,000)=240
560-320=240
240=240